# Don’t Retrieve, Navigate: Distilling Enterprise Knowledge into Navigable Agent Skills for QA and RAG

### 29 Apr 2026 v2

@@ Appendix B Skill Structure Examples
We show the full content of representative navigation files from the WixQA compilation (default configuration: p=10p{=}10, K=7K{=}7). These files are what the agent reads during navigation.
@@
This file sits at the top of the skill-01-wix-commerce-monetization skill directory, the largest skill covering 1,513 documents across 16 subgroups. The agent reads this file in a single tool call and uses the subgroup summaries to decide which branch to explore.
name: skill-01-wix-commerce-monetization
@@
  This group covers the full Wix commerce
  and monetization ecosystem, spanning every
  major product vertical and workflow
  available to Wix merchants and business
  owners. Topics range from storefront and
  product management (Wix Stores,
  Restaurants, POS, Print/Fulfillment) to
  financial operations (payments, taxes,
  refunds, invoices, pricing plans,
  subscriptions, and promotional tools like
  coupons and gift cards)...
@@
This group covers the full Wix commerce and
monetization ecosystem, spanning every major
product vertical and workflow. Topics range
from storefront and product management (Wix
Stores, Restaurants, POS, Print/Fulfillment)
to financial operations (payments, taxes,
refunds, invoices, pricing plans,
subscriptions, and promotional tools), as
well as loyalty programs, partner ecosystems,
and learning/course delivery via Wix Online
Programs. Users can get answers about setting
up and running an online or in-person
business on Wix.
@@
- **group-00-wix-restaurants-platform/**
  (159 docs): Covers the Wix Restaurants
  platform in its entirety, spanning both
  the original and newer versions of Wix
  Restaurants Orders and Menus apps.
- **group-01-wix-payments-ecosystem/**
  (87 docs): Covers Wix’s payment ecosystem,
  spanning payment processing, account
  management, and financial operations.
- **group-02-wix-pos-systems/** (110 docs):
  Covers Wix’s Point of Sale (POS)
  ecosystem, both tablet-based Retail POS
  and smartphone-based Mobile POS.
- **group-03-wix-stores-product-management/**
  (99 docs): Covers Wix Stores product and
  storefront management, from setup to
  customization.
- **group-04-wix-stores-shipping-config/**
  (57 docs): Covers shipping, pricing, and
  checkout configuration.
- **group-05-wix-payment-providers/**
  (289 docs): Covers payment processing and
  provider integration, both first-party
  and third-party gateways.
- **group-06-wix-subscription-management/**
  (133 docs): Covers subscription and pricing
  plan management across Premium, Studio,
  Editor X, Ascend, and App Market plans.
- **group-07-wix-print-fulfillment/**
  (55 docs): Covers print and fulfillment-
  related commerce features.
- **group-08-wix-commerce-fulfillment/**
  (74 docs): Covers order management,
  fulfillment, shipping, and inventory.
- **group-09-wix-online-programs/**
  (63 docs): Covers Wix Online Programs for
  building and delivering online courses.
- **group-10-refund-cancellation-policies/**
  (25 docs): Covers refund and cancellation
  policies across Wix products.
- **group-11-wix-pricing-plans/**
  (113 docs): Covers Wix Pricing Plans and
  monetization features lifecycle.
- **group-12-wix-partner-loyalty-programs/**
  (75 docs): Covers partner, loyalty, and
  rewards ecosystems.
- **group-13-taxes-fees-config/**
  (53 docs): Covers tax configuration,
  tips, and fees/surcharges.
- **group-14-wix-billing-documents/**
  (87 docs): Covers financial document
  management: invoices, quotes, proposals,
  receipts.
- **group-15-wix-payments-promotions/**
  (34 docs): Covers promotional tools:
  discount coupons, gift cards, and saved
  payment methods.
@@
This file sits one level below the root, inside group-01-wix-payments-ecosystem/, covering 87 documents across 10 leaf groups. The agent reads this after deciding from the SKILL.md above that payments are relevant.
name: wix-payments-ecosystem
@@
  Covers Wix’s payment ecosystem, spanning
  payment processing, account management,
  and financial operations across Wix
  Payments and related commerce tools.
@@
These sub-groups collectively cover Wix’s
payment ecosystem, spanning payment
processing, account management, and
financial operations across Wix Payments
and related commerce tools (Wix Stores,
Bookings, Restaurants, and Forms). Topics
range from setting up and verifying Wix
Payments accounts, managing payouts and
transactions, handling fraud, chargebacks,
and restricted products, to workflows like
collecting donations, processing manual
card entries, and understanding split
payment limitations.
@@
Read the INDEX.md in each sub-group to
understand what it covers.
- **group-00/** (6 docs): Wix account
  billing, payments management, and
  role-based permissions.
- **group-01/** (14 docs): Wix Payments
  payment processing, payout management,
  and account administration.
- **group-02/** (15 docs): Setup, mgmt,
  and compliance requirements of Wix
  Payments accounts.
- **group-03/** (7 docs): Accepting and
  managing online donations through Wix.
- **group-04/** (8 docs): Payment
  flexibility limitations across Wix
  products (split, defer, partial pay).
- **group-05/** (5 docs): Fraud prevention
  and payment security within Wix Payments.
- **group-06/** (8 docs): Wix Payments
  dashboard and transaction management.
- **group-07/** (12 docs): Credit and debit
  card payment processing within Wix.
- **group-08/** (4 docs): Payment provider
  compliance and restricted/prohibited
  products.
- **group-09/** (8 docs): Chargebacks and
  fees within Wix Payments.
@@
This file sits at the bottom of the hierarchy, inside group-05/ under the payments ecosystem. It lists individual document IDs with titles. The agent reads this to select which documents to retrieve via get_document.
name: L1-C298
@@
  Fraud prevention and payment security
  within Wix Payments, addressing merchant-
  facing and customer-facing concerns
  related to credit card transactions.
@@
These documents cover fraud prevention and
payment security within Wix Payments,
addressing both merchant-facing and
customer-facing concerns related to credit
card transactions. They answer questions
about how to identify and respond to
suspicious or declined payments, understand
liability for fraudulent chargebacks, and
interpret unusual bank statement activity.
Key terms: liability shift, chargeback
mechanisms, chip card readers, suspicious
payment flags, fraud detection models.
@@
### Documents (5 items)
Use get_document tool with the doc_id
to read full content.
- ‘56ae7642b2a8c2ea‘: Best Practices for
  Accepting Card Payments
- ‘448d7947c5a22489‘: Wix Payments:
  About the ’Liability Shift’
- ‘bb73d9cb90385f12‘: Wix Payments:
  Declined Payment Due to Suspected Fraud
- ‘c3b01aac19482333‘: Wix Payments:
  About Suspicious Payments
- ‘d22b8cf6e599e0b8‘: Double Charge
  Appearing on Your Bank Statement
@@ Appendix D Serving System Prompt
into topic clusters.
@@
- SKILL.md files are NAVIGATION AIDS --
  they tell you where to look, not what
  to say.
@@
## Navigation Strategy
Your skills directory has this structure:
@@
    SKILL.md        <- top-level summary
                       + index of contents
@@
      INDEX.md      <- summary + index of
                       sub-contents
@@
        INDEX.md    <- summary + document
                       ID listings
Skill names and descriptions are already
available to you. Follow this workflow:
1. Read the SKILL.md of the 1-2 most
   relevant skills for your query.
2. Drill into the most relevant sub-group:
   Read its INDEX.md.
3. At the leaf level, INDEX.md lists
   document IDs with brief titles. Pick
   the most relevant document IDs.
4. Call get_document with each relevant
   doc_id to retrieve the full text.
5. Read at least one full document before
   answering.
@@ 5. Read at least one full document before
- Code execution: Use ‘ls‘ and ‘cat‘ to
  navigate the skills hierarchy.
@@
  text of a document by its ID. The doc_id
  values are listed in leaf-level INDEX.md
  files.
@@ Appendix E Full Navigation Traces
"wix-commerce-monetization" covers payments,
@@
Tool: text_editor_code_execution
  view skill-01-wix-commerce-monetization/
@@
Result: 10 subgroups listed, including:
  - group-01-wix-payments-ecosystem/
    (87 docs): "payment processing,
    account management, financial
    operations"
@@
Tool: text_editor_code_execution
  view group-01-wix-payments-ecosystem/
    INDEX.md
Result: 8 leaf groups. Agent identifies
  group-05 covering Wix Payments account
  setup and verification.
Tool: text_editor_code_execution
  view group-05/INDEX.md
Result: 12 documents listed, including:
  - d56cc797742e: "Changing Your Wix
    Payments Account Type"
--- Turn 3 ---
@@
Navigation: 6221 -> 1513 -> 87 -> 12 -> 1
Trace 2: Multi-document retrieval (Factuality 0.8, CtxR 1.0).
@@
Agent thought: Hotels and scheduling fall
under "wix-business-tools" which covers
scheduling, CRM, and business management.
Tool: text_editor_code_execution
  view skill-00-wix-business-tools/SKILL.md
Result: 8 subgroups listed, including:
  - group-02-wix-hotels-platform/
    (82 docs): "hotel management, bookings,
    reservations, calendar sync"
--- Turn 2 ---
Tool: text_editor_code_execution
  view group-02-wix-hotels-platform/INDEX.md
Result: 7 leaf groups. Agent identifies
  group-05 covering external calendar
  integration and iCal sync.
Tool: text_editor_code_execution
  view group-05/INDEX.md
@@
under "wix-platform-guide" which covers
Editor, site building, and customization.
Tool: text_editor_code_execution
  view skill-03-wix-platform-guide/SKILL.md
Result: 10 subgroups. Agent picks
  "wix-mobile-site-building" covering
@@
--- Turn 2 ---
Tool: text_editor_code_execution
  view wix-mobile-site-building/INDEX.md
@@
  elements. Agent finds group with docs
  on Quick Action Bar.
Tool: text_editor_code_execution
  view group-03/INDEX.md
Result: 8 documents including:
@@
Navigation: 6221 -> 1161 -> 142 -> 8 -> 2
Trace 4: Cross-branch exploration (Factuality 1.0, CtxR 0.6).
@@
Programs, and "currency" is a payments/
billing topic. The skill "wix-commerce-
monetization" covers both online programs
and payment configuration.
Tool: text_editor_code_execution
  view skill-01-wix-commerce-monetization/
@@
Result: 15 subgroups listed, including:
  - group-09-wix-online-programs/
    (107 docs): "online courses, program
    management, pricing, delivery"
  - group-14-wix-billing-documents/
    (68 docs): "currency settings, invoice
    automation, financial documents"
--- Turn 2 ---
Agent thought: Start with Online Programs
since the query is about course currency.
Tool: text_editor_code_execution
  view group-09-wix-online-programs/INDEX.md
Result: 5 leaf groups. Picks group-03
  covering program settings (18 docs).
Tool: text_editor_code_execution
  view group-03/INDEX.md
Result: 18 documents including:
  - 7a9dc6f2de90: "Online Programs:
    Editing Your Program Settings"
--- Turn 3 ---
@@
  {"doc_id": "7a9dc6f2de90db61"}
Result: Program settings article. Mentions
@@
configuration steps. Navigate to billing.
Tool: text_editor_code_execution
  view group-14-wix-billing-documents/
    INDEX.md
Result: 5 leaf groups. Picks group-00
  covering currency management (18 docs).
Tool: text_editor_code_execution
  view group-00/INDEX.md
Result: 18 documents including:

### 15 May 2026 v3

@@ Appendix B Skill Structure Examples
We show the full content of representative navigation files from the WixQA compilation (default configuration: p=10p{=}10, K=10K{=}10). These files are what the agent reads during navigation.
@@
This file sits at the top of the skill-00-wix-commerce-operations skill directory, the largest skill covering 1,682 documents across 11 subgroups. The agent reads this file in a single tool call and uses the subgroup summaries to decide which branch to explore.
name: skill-00-wix-commerce-operations
@@
  This group covers the Wix commerce and
  business operations ecosystem, spanning
  the full range of tools for running
  transactions, managing services, and
  scaling a business on Wix. Topics include
  e-commerce store management, restaurant
  and hotel operations, event ticketing,
  point-of-sale systems, bookings and
  scheduling, pricing plans, payments and
  payment providers, tax and invoicing,
  promotions and discounts...
@@
This group covers the Wix commerce and
business operations ecosystem, spanning
the full range of tools for running
transactions, managing services, and
scaling a business on the Wix platform.
Topics include e-commerce, restaurant and
hotel operations, event ticketing, POS,
bookings, pricing plans, payments, tax,
invoicing, promotions, and multilingual
localization. Users can get answers about
setting up and configuring any Wix commerce
product and processing payments across
global providers.
@@
- **group-00-wix-events-platform/**
  (138 docs): Wix Events platform: event
  creation, ticketing, registration, guest
  management, and promotion.
- **group-01-wix-restaurants-platform/**
  (232 docs): Wix Restaurants: menus,
  ordering, reservations, delivery, POS
  integrations, and fulfillment.
- **group-02-wix-hotels-bookings/**
  (109 docs): Wix Hotels and Bookings:
  room configuration, pricing, calendar
  sync, and OTA channel management.
- **group-03-wix-pos-ecosystem/**
  (167 docs): Wix POS ecosystem: hardware
  setup, payment processing, staff
  management, and inventory.
- **group-04-wix-payments-processing/**
  (113 docs): Payments lifecycle: account
  setup, chargebacks, fraud prevention,
  payouts, and subscription billing.
- **group-05-wix-billing-tax-management/**
  (110 docs): Financial and tax management:
  invoicing, quotes, Avalara, VAT/GST.
[... 5 more sub-groups omitted ...]
@@
This file sits one level below the root, inside group-04-wix-payments-processing/, covering 113 documents across 10 leaf groups. The agent reads this after deciding from the SKILL.md above that payments are relevant.
name: wix-payments-processing
@@
  Covers the broad domain of payments on
  the Wix platform, spanning the full
  lifecycle from account setup to
  transaction management, payouts, and
  compliance.
@@
Topics range from setting up and verifying
Wix Payments accounts, enabling third-party
providers (PayPal, Stripe, Apple Pay,
Google Pay), and managing chargebacks and
fraud prevention, to payout schedules,
prohibited product restrictions,
subscription billing, and unsupported
payment features.
@@
- **group-00-wix-payments-chargebacks/**
  (12 docs): Chargeback management,
  transaction fees, and merchant practices.
- **group-01-wix-payment-setup/**
  (22 docs): Setup and troubleshooting of
  payment methods across commerce products.
- **group-02-wix-payments-payouts/**
  (15 docs): Payout management, settlement
  reports, and fund transfer timelines.
- **group-03-payment-flexibility-requests/**
  (9 docs): Feature requests for split,
  installment, and deferred payments.
- **group-04-wix-paypal-button-setup/**
  (7 docs): PayPal Buy Now and Donate
  button configuration.
- **group-05-wix-payments-account-setup/**
  (13 docs): Account setup, verification,
  and EIN/identity documents.
- **group-06-wix-payments-processing/**
  (10 docs): Payment processing issues,
  fraud prevention, and declined payments.
- **group-07-wix-payments-management/**
  (9 docs): Transaction dashboards, payment
  history, and settlement reports.
- **group-08-prohibited-payment-products/**
  (6 docs): Payment provider compliance and
  restricted product categories.
- **group-09-wix-subscription-billing/**
  (10 docs): Premium plan billing, renewal
  payments, and payment method updates.
@@
This file sits at the bottom of the hierarchy, inside group-06-wix-payments-processing/ under the payments processing branch. It lists individual document IDs with titles and summaries. The agent reads this to select which documents to retrieve via get_document.
name: wix-payments-processing
@@
  Payment processing issues, fraud
  prevention, and transaction management
  within Wix Payments.
@@
These documents cover payment processing
issues, fraud prevention, and transaction
management within Wix Payments. They answer
questions about why payments are declined
(AVS mismatches, CVV errors, 3-D Secure
failures, suspected fraud), how to resolve
double charges, and how EMV chip liability
shift rules affect merchants.
@@
### Documents (10 items)
Use get_document with the doc_id to
read full content.
- ‘56ae7642b2a8c2ea‘ -- Best Practices
  for Accepting Card Payments
- ‘88f6b5848f6f662d‘ -- Troubleshooting
  Payment Failures and Declined
  Transactions
- ‘d22b8cf6e599e0b8‘ -- Double Charge on
  Bank Statement Explanation
- ‘448d7947c5a22489‘ -- Wix Payments:
  Liability Shift and Chip Card
  Requirements
- ‘bb73d9cb90385f12‘ -- Wix Payments
  Declined Payment Suspected Fraud
[... 5 more documents omitted ...]
@@ 3. What types of user QUESTIONS this group
Document summary card.
Issued once per document at the start of compilation. The returned JSON card is concatenated with the truncated raw text for embedding and is reused as INDEX.md row metadata.
You are summarizing a single document from
a knowledge base. Produce a compact JSON
’card’ describing the document so an agent
can scan many cards quickly and pick the
relevant one.
Return ONLY a JSON object with exactly
these keys:
  title:    short descriptive title
            (5-12 words). Reuse the
            document’s own title if it has
            one, otherwise infer.
  one_line: a single sentence (max 25
            words) describing what the
            document covers and what kind
            of question it answers.
  phrases:  list of 2-4 short distinctive
            phrases (2-6 words each) that
            are characteristic of this
            document and unlikely to appear
            in unrelated documents (e.g.
            product names, schema names,
            specific entities, code
            identifiers, section headers).
Be specific and concrete. Do not include
generic phrases like ’knowledge base
article’ or ’support document’.
Document:
{document}
JSON card:
Verify-and-repartition.
Issued once per hierarchy level (skipped beyond 60 sibling clusters per call). The model is asked to flag confusable cluster pairs and mixed individual clusters, and to return ID-level reassignments. Changes are applied verbatim; affected clusters are re-summarized.
You are auditing a clustering of items at
level {level} of a navigable knowledge
hierarchy. Each cluster has a label, a
short summary, and a list of member item
IDs with brief titles.
Two failure modes to look for:
  (a) CONFUSABLE: two or more clusters
      whose labels/summaries describe
      essentially the same topic.
  (b) MIXED: a single cluster whose
      members visibly belong to multiple
      distinct sub-topics that should be
      split.
For each problematic GROUP (a set of
confusable cluster IDs, or a single mixed
cluster ID), output a fresh partition of
its members by ID. Use ONLY the IDs already
listed in that group; do not invent IDs or
move items across non-flagged boundaries.
Return ONLY a JSON object with this shape:
{ "changes": [
  { "old_cluster_ids": ["L1-C2", "L1-C7"],
    "new_partition": {
      "label_a": ["id1","id2"],
      "label_b": ["id3","id4"]
If no changes are needed, return
{ "changes": [] }.
Clusters at level {level}:
{clusters_block}
JSON:
Entity and doc-type extraction.
Issued once per cluster, batched concurrently, after the hierarchy stabilizes. The returned named entities are aggregated into entity_index.json and drive the per-skill Related skills cross-references.
You are tagging a topic cluster from a
knowledge base with the entities and
document types it contains.
Extract ONLY two categories:
  named_entities: proper-noun product
    names, feature names, person names,
    place names, organization names,
    schema names, specific identifiers.
    Exclude generic concepts (e.g.
    ’authentication’, ’payments’ are
    generic; ’Stripe Connect’,
    ’Wix Bookings’ are named entities).
  doc_types: kinds of documents present
    in the cluster, e.g. ’FAQ’, ’release
    notes’, ’balance sheet’, ’contract
    clause’, ’API reference’, ’tutorial’,
    ’changelog’.
Return ONLY a JSON object:
{ "named_entities": ["..."],
  "doc_types": ["..."] }
Aim for 3-10 named entities and 1-3 doc
types. Use lowercase, no duplicates.
Cluster label: {label}
Cluster summary: {summary}
Example item titles:
{titles_block}
JSON:
@@ Appendix D Serving System Prompt
into topic clusters, and you have access
to an entity cross-index and a leaf-level
search tool to triangulate evidence.
@@
- SKILL.md / INDEX.md files are NAVIGATION
  AIDS -- they tell you where to look,
  not what to say.
@@
## Skill directory layout
@@
  SKILL.md          <- top-level summary,
                       exemplar docs,
                       related skills,
                       index
@@
    INDEX.md        <- sub-group summary
                       + child index
@@
      INDEX.md      <- leaf summary +
                       document rows
                       (id + title +
                       phrases)
Each SKILL.md / INDEX.md may contain:
- ## Overview      -- what this skill
                       covers
- ## Related skills -- sibling skills that
                       share entities;
                       cross-jump here when
                       the current branch
                       is thin
- ## Entities & document types
                    -- quick scan of named
                       entities present
- ## Example documents in this skill
                    -- representative items
                       (read one of these
                       to learn what kind
                       of content lives
                       here)
- ## Contents       -- sub-groups and/or
                       document rows
- ### See also      -- @see stubs
                       (documents whose
                       primary entry is
                       elsewhere)
A corpus-level entity_index.json is also
present at the corpus root. It maps each
named entity to the skill paths that
mention it; use it to find ALL skills
relevant to a query entity rather than
only the first match.
## Navigation Strategy (multi-candidate)
1. SCAN at least 2 candidate top-level
   skills before committing to a path.
   - If the query mentions a specific
     entity, first
     ‘cat entity_index.json | grep -i
       <entity>‘
     to find every skill that references
     it, and read the SKILL.md of the top
     2 by mention count.
   - Otherwise read the SKILL.md of the 2
     most-plausible top-level skills.
2. For each candidate, descend ONE level
   (read its first relevant INDEX.md)
   before pruning. Write down which
   candidates you keep and why.
3. At the leaf, use ‘cat ... | grep ...‘
   against the INDEX.md doc rows to pick
   the most promising doc_ids -- the rows
   include the document title, one-liner,
   and key phrases, which is usually
   enough to triangulate without opening
   anything.
4. If two candidate paths return
   comparable evidence, retrieve documents
   from BOTH and explicitly compare. Pick
   the best answer; do NOT just take the
   first match.
5. If a candidate skill has a ‘## Related
   skills‘ block and your initial path is
   thin, cross-jump to a related skill
   before giving up.
6. Read at least one full document via
   get_document before answering.
@@ 6. Read at least one full document via
- Code execution: Use ‘ls‘, ‘cat‘,
  ‘cat ... | grep ...‘ to navigate the
  skills hierarchy and the
  entity_index.json.
@@
  text of a document by its ID.
@@ Appendix E Full Navigation Traces
"wix-commerce-operations" covers payments,
@@
Tool: code_execution
  cat skill-00-wix-commerce-operations/
@@
Result: 11 subgroups listed, including:
  - group-04-wix-payments-processing/
    (113 docs): "payments lifecycle: account
    setup, chargebacks, fraud prevention,
    payouts, and subscription billing"
Tool: code_execution
  cat group-04-wix-payments-processing/
    INDEX.md
Result: 10 leaf groups. Agent identifies
  group-05-wix-payments-account-setup
  covering account setup and verification.
Tool: code_execution
  cat group-05-wix-payments-account-setup/
    INDEX.md
Result: 13 documents listed, including:
  - d56cc797742e: "Switching Between
    Individual and Business Account Types
    in Wix Payments"
@@
Navigation: 6221 -> 1682 -> 113 -> 13 -> 1
Trace 2: Multi-document retrieval (Factuality 1.0, CtxR 1.0).
@@
Agent thought: Hotels and calendar sync
fall under "wix-commerce-operations" which
covers hotel operations and bookings.
Tool: code_execution
  cat skill-00-wix-commerce-operations/
    SKILL.md
Result: 11 subgroups listed, including:
  - group-02-wix-hotels-bookings/
    (109 docs): "Wix Hotels and Bookings:
    room configuration, pricing, calendar
    sync, and OTA channel management"
Tool: code_execution
  cat group-02-wix-hotels-bookings/
    INDEX.md
Result: 8 leaf groups. Agent identifies
  group-07-hotel-channel-manager-setup
  covering external calendar integration.
Tool: code_execution
  cat group-07-hotel-channel-manager-setup/
    INDEX.md
@@
under "wix-platform-overview" which covers
editors, CMS, and design.
Tool: code_execution
  cat skill-01-wix-platform-overview/
    SKILL.md
Result: 14 subgroups. Agent picks
  group-13-wix-mobile-apps covering
@@
Tool: code_execution
  cat group-13-wix-mobile-apps/INDEX.md
@@
  elements. Agent finds group-00 with docs
  on Quick Action Bar (6 docs).
Tool: code_execution
  cat group-00-mobile-quick-action-bar/
    INDEX.md
Result: 6 documents including:
@@
Navigation: 6221 -> 1770 -> 60 -> 6 -> 1
Trace 4: Cross-branch exploration (Factuality 1.0, CtxR 1.0).
@@
Programs under "wix-platform-ecosystem",
and "currency" is a payments topic under
"wix-commerce-operations".
Tool: code_execution
  cat skill-02-wix-platform-ecosystem/
@@
Result: 11 subgroups listed, including:
  - group-07-wix-programs-partnerships/
    (119 docs): "online programs, courses,
    pricing, and partner integrations"
Tool: code_execution
  cat group-07-wix-programs-partnerships/
    INDEX.md
Result: 6 leaf groups. Picks group-04
  covering online program setup (9 docs).
Tool: code_execution
  cat group-04-wix-online-programs-setup/
    INDEX.md
Result: 9 documents including:
  - 3b2e9b0c4273: "Wix Online Programs:
    Setting Course Pricing"
@@
  {"doc_id": "3b2e9b0c4273f8c9"}
Result: Course pricing article. Mentions
@@
configuration steps. Navigate to payments.
Tool: code_execution
  cat skill-00-wix-commerce-operations/
    SKILL.md | grep -i currency
Result: group-09-payment-provider-
  integration mentions currency settings.
Tool: code_execution
  cat group-09-payment-provider-integration/
    group-03-wix-currency-settings/INDEX.md
Result: 12 documents including: